OpenStudy (anonymous):

logarithmic form of 49=(1/7)^-2

OpenStudy (nurali):

49=(1/7)^-2 49=1/(1/7)^2 49=1/1/49 49=49

OpenStudy (debbieg):

\[\log_{1/7} 49=-2\]